A THEORETICAL STUDY OF NiCN IN THE ^{2}Δ ELECTRONIC GROUND STATE

\underline{TSUNEO HIRANO}, REI OKUDA, and UMPEI NAGASHIMA, Research Institute for Computational Sciences, National Institute of Advanced Industrial Science and Technology, 1-1-1 Umezono, Tsukuba, Ibaraki 305-8568, Japan; PER JENSEN, Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

The three-dimensional potential energy surface of X̃^{2}Δ_{i} NiCN has been calculated ab initio at the MR-SDCI+Q+E_{rel}/[Roos ANO (Ni), aug-cc-pVQZ (C, N)] level of theory. The equilibrium geometry derived from this surface is linear with r_{e}(Ni-C) = 1.814 [1.8292(28), 1.8293(1)] Å and r_{e}(C-N) = 1.167 [1.1591(29), 1.1590(2)] Å, where the values in brackets are r_{0} values for the ground Ω = 5/2 spin-substate determined experimentally by Kingston et al.^{a} and Sheridan et al.,^{b} respectively. From the electronic structure given in terms of natural orbitals, and the Mulliken population^{c} of +0.83 on Ni, we conclude that the Ni-C bond is basically ionic but less ionic than those of FeNC and CoCN. The electron from Ni goes into the Ni-mediated CN σ* orbital, giving the electron distribution Ni^{+0.8}(CN)^{−0.8}. The 3d-π* backbonding is not observed. Molecular constants determined from the ab initio potential energy surface by perturbation methods and in variational calculations will be reported: For example, ω_{1} = 2198 cm^{−1}, ω_{2} = 254 cm^{−1}, and ω_{3} = 511 cm^{−1}. A severe Fermi resonance between 2ν_{2} and ν_{3} is expected. A spin-orbit interaction scheme including the ab initio predicted spin-orbit coupling constant A_{SO} = −613 cm^{−1}^{d} will be presented.

aC. T. Kingston, A. J. Merer, and T. D. Varberg, J. Mol. Spectrosc., 215, 106 (2002).

bP. M. Sheridan and L. M. Ziurys, J. Chem. Phys., 118, 6370 (2003).

cComputed at the MR-SDCI/[Wachters+f (Ni), aug-cc-pVTZ (C, N)] level of theory

dcf. the unperturbed A_{SO}-value of −594.2(5) cm^{−1} for X^{2}Δ NiH; J. A. Gray, M. Li, T. Nelis, and R. W. Field, J. Chem. Phys., 95, 7164 (1991).